# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v3

@@ 5 Analysis
To isolate the impact of initialization, we take truncated SVD on various matrices, including Kaiming initialization (He et al., 2015) and $\Delta W_{avg}$ with varying levels of Gaussian noise, as shown in Table 4. By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. This results in $B^{T}B=AA^{T}=I$, allowing us to isolate the effect of initialization. The results clearly demonstrate the significance of initialization—our approach consistently outperforms other variants.
@@
$B_{\text{init}}\leftarrow U[1:r]S[1:r,1:r]$ (15)
$A_{\text{init}}\leftarrow V[1:r]$ (16)
$R_{\text{init}}\leftarrow I$ (17)
This construction ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$. The setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3. We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B on MetaMathQA, as shown in Figure 3. Although both start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation. Furthermore, LoRA-SB achieves higher accuracies on GSM8K and MATH, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
Figure 3: Training loss for Mistral-7B on MetaMathQA, highlighting the impact of optimal gradient approximation.
Training Time Overhead vs LoRA-XS.
As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset. Table 5 presents the associated training time overhead for these computations, compared to LoRA-XS. The results show that the additional overhead is negligible, adding just 24242–42 – 4 minutes compared to the total training time of 35353–53 – 5 hours per epoch ($\approx 1.1\%$ to $1.3\%$). Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
Model Overhead Training Time/Epoch
Mistral 7B 0:02:01 3:03:57
Gemma-2 9B 0:03:46 4:13:24
Llama-3.2 3B 0:03:54 4:54:31
Inference Overhead vs LoRA
LoRA-SB introduces a minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, and the need for higher ranks to achieve comparable performance to LoRA. We benchmark the inference-time FLOPs and MACs across various models and find that the overhead is negligible. A detailed analysis can be found in Table 6 of Appendix B.
@@ 6 Conclusion
In this work, we introduced LoRA-SB, which bridges the gap between low-rank PEFT and full FT. This is enabled by our initialization strategy, which approximates the first step of full FT and ensures that the most relevant subspaces for task-specific adaptation are captured. We achieve optimal gradient scaling and preserve update directions throughout training. Our approach ensures hyperparameter independence by approximating the full FT gradient, thereby eliminating instability issues associated with scaling factors.
Through extensive experiments, we demonstrate that our method outperforms LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS. Our work advances PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks. Future work includes exploring adaptive layer-wise rank settings and integrating LoRA-SB with quantization. We also aim to evaluate its performance on other models, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
@@ 7 Acknowledgements
Appendix A Mathematical Proofs
@@ A.1 Proof of Lemma 1
Let $\Delta W$ be an update learned with LoRA-XS. Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}\mid\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where for any $M\in\mathbb{R}^{m\times n}$ the column and row spaces are defined as $\text{Col}(M)=\{y\in\mathbb{R}^{m}\mid\exists x\in\mathbb{R}^{n}:\;y=Mx\}$ and $\text{Row}(M)=\{y\in\mathbb{R}^{n}\mid\exists x\in\mathbb{R}^{m}:\;y=M^{\top}x\}$.
@@
$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\subseteq\text{Col}(B).$
@@ A.2 Proof of Lemma 2
The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
@@
$y\frac{\partial X}{\partial R}=yA^{\top}\implies\frac{\partial L}{\partial R}=yA^{\top}=sB^{\top}gA^{\top}$ (22)
@@
A.3 Proof of Theorem 3
Theorem.
The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (24)
@@ A.3 Proof of Theorem 3
[…]
A.4 Proof of Theorem 4
@@
Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate. This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0$
@@
Next, we note that matrices $B^{\top}B\in\mathbb{R}^{r\times r}$ and $AA^{\top}\in\mathbb{R}^{r\times r}$ are positive definite since they are positive semi-definite and matrices $B$ and $A$ are full-rank (i.e., with rank $r$) matrices, which means that $B^{\top}B$ and $AA^{\top}$ have non-zero eigenvalues. Therefore, $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ are also positive definite, implying that there exist matrices $U$ and $V$ such that $(B^{\top}B)^{-1}=VV^{\top}$ and $(AA^{\top})^{-1}=UU^{\top}$ (e.g., one can find such matrices using Cholesky decomposition). Then, we have
$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}$ $=\langle g^{R}_{\text{LoRA-XS}},VV^{\top}g^{R}_{\text{LoRA-XS}}UU^{\top}\rangle_{F}$
$=\frac{1}{s^{2}}\langle V^{\top}g^{R}_{\text{LoRA-XS}}U,V^{\top}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
$=\|V^{\top}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0.$
@@ A.5 Proof of Theorem 5
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\quad\text{ but not when }\quad\tilde{g}=sBg^{R}_{LoRA-XS}A$
@@ A.6 Proof of Theorem 6
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
@@
Appendix B Inference Overhead vs LoRA
As discussed in Section 5, LoRA-SB introduces minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, as well as the requirement for higher ranks to match LoRA’s performance. This comparison is presented in Table 6, showing that the additional overhead of LoRA-SB is negligible.
@@ Appendix B Inference Overhead vs LoRA
RoBERTa-large LoRA $8$ $77.86$ G $155.79$ G
@@
Appendix C Experiment Details
We ran our experiments on a single NVIDIA A6000 GPU, averaging results over three independent trials. We trained all models using the AdamW optimizer (Loshchilov & Hutter, 2019).
For arithmetic and commonsense reasoning tasks, we set up Mistral-7B, Gemma-2 9B, and Llama-3.2 3B with hyperparameters and configurations listed in Table 7. We adopted most settings from previous studies (Hu et al., 2023) but conducted our own learning rate sweep. Following LoRA-XS guidelines, we set $\alpha=r$ for their baseline configuration.
For the GLUE benchmark using RoBERTa-large, you can find the hyperparameter details in Table 8. We mostly adhered to the original configurations from the LoRA paper (Hu et al., 2021) but adjusted the learning rate through a sweep. In line with LoRA-XS settings, we fixed $\alpha$ at $16$ for their baseline.
For all tasks, we followed the baseline configurations provided in the PiSSA (Meng et al., 2024) and rsLoRA (Kalajdzievski, 2023) papers for our comparisons.
@@ Appendix C Experiment Details
Appendix D Dataset Details
The MetaMathQA dataset (Yu et al., 2024) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information. We assess this dataset using two benchmarks: GSM8K (Cobbe et al., 2021), which consists of grade-school math problems requiring multi-step reasoning, and MATH (Hendrycks et al., 2021), which presents difficult, competition-level math problems. Evaluation focuses solely on the final numeric answer.
CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023). Each example is framed as a multiple-choice question where the model generates the correct answer without explanations. We use the prompt template from (Hu et al., 2023). The individual datasets used are described below:

### v4

@@ 4 Analysis
To isolate the impact of initialization, we take truncated SVD on various matrices, including Kaiming initialization (15) and $\Delta W_{avg}$ with varying levels of Gaussian noise, as shown in Table 4. By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. This results in $B^{T}B=AA^{T}=I$, allowing us to isolate the effect of initialization. The results clearly demonstrate the significance of initialization, our approach consistently outperforms other variants.
@@
Why Do We Use 0.1% of the Dataset Size for Initialization?
We selected the $0.1\%$ initialization dataset-size heuristic based on experiments that suggested it provides a good tradeoff between quality and efficiency. Specifically, we conducted ablations varying the number of samples used for initialization when fine-tuning Mistral-7B and Gemma-2 9B on 50k samples from MetaMathQA. The results (Table 5) show that once the sample count exceeds a modest threshold (25 samples or $0.05\%$), performance quickly plateaus, indicating that the learned subspace is already sufficiently representative. Using $0.1\%$ of the training data (50 samples) consistently exceeds this threshold across tasks and models, while incurring negligible training time overhead.
Table 5: Performance effect of number of samples used for initialization.
# Samples Mistral-7B Gemma-2 9B
GSM8K ($\uparrow$) MATH ($\uparrow$) GSM8K ($\uparrow$) MATH ($\uparrow$)
11 $62.13$ $15.55$ $76.03$ $35.77$
55 $62.78$ $16.86$ $77.49$ $37.24$
$25$ $63.28$ $17.30$ $78.18$ $37.70$
$50$ $63.38$ $17.44$ $78.40$ $37.70$
$100$ $63.34$ $17.25$ $78.22$ $37.45$
$200$ $63.45$ $17.36$ $78.43$ $37.87$
$500$ $63.40$ $17.52$ $78.54$ $37.63$
@@
$B_{\text{init}}\leftarrow U[1:r]S[1:r,1:r],A_{\text{init}}\leftarrow V[1:r],R_{\text{init}}\leftarrow I$ (12)
This ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$. The setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3. We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B, as shown in Figure 3 in Appendix E. Although both start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation. Furthermore, LoRA-SB achieves higher accuracies on GSM8K and MATH, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
Training Time and Inference.
We provide detailed benchmarks of training time and inference performance in Appendix F and G, respectively. As shown, the initialization step in LoRA-SB introduces only a negligible training-time overhead compared to LoRA ($\approx 1.1\%-1.3\%$).
@@ 5 Conclusion
In this work, we introduced LoRA-SB, which bridges the gap between low-rank PEFT and full FT. This is enabled by our initialization strategy, which approximates the first step of full FT and ensures that the most relevant subspaces for task-specific adaptation are captured. We achieve optimal gradient scaling and preserve update directions throughout training. Our approach ensures scaling factor independence by approximating the full FT gradient, thereby eliminating potential instability issues. Through extensive experiments, we demonstrate that our method outperforms LoRA (and baselines) using upto 90x less parameters, and comprehensively outperforms LoRA-XS.
@@ 6 Acknowledgements
Appendix
Appendix A Related Work
Parameter-Efficient Fine-Tuning (PEFT). PEFT methods have become essential for adapting large pre-trained models under computational constraints. Early techniques like AdapterFusion (34) and Prefix-Tuning (24) enabled task-specific adaptation with minimal parameter updates. Advances like soft prompts (23) further reduced trainable parameter counts while maintaining strong performance. Recent approaches have explored operating directly on model representations (49).
Low-Rank Decomposition Methods. LoRA (17) demonstrated that weight updates during FT could be efficiently approximated using low-rank matrices, drastically reducing parameter counts. Building on this insight, variants such as QLoRA (9) and AdaLoRA (52) extended the paradigm through quantization and adaptive allocation strategies. The applicability of low-rank techniques has also been explored in pretraining with GaLore (53) and ReLoRA (25), highlighting the versatility of low-rank adaptation methods. LoRA-based methods have also been applied in other domains, such as efficient federated FT (42, 40).
Enhancing LoRA Performance. Recent efforts have focused on optimizing LoRA’s performance. PiSSA (30) demonstrated improvements by initializing matrices with principal components of pre-trained weights. LoRA-Pro (46) and LoRA-GA (45) improved gradient approximation, aligning low-rank updates more closely with full FT. Methods like DoRA (26) and rsLoRA (20) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
Improving Efficiency in LoRA Variants. Efficiency-focused innovations have pushed LoRA toward more parameter savings. LoRA-XS (2) achieves this by inserting a small trainable weight matrix into frozen low-rank matrices. VeRA (22) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation. Tied-LoRA (37) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (44) introduces an asymmetric architecture for improvement.
Appendix B Proofs
@@ B.1 Proof of Lemma 1
Let $\Delta W$ be an update learned with LoRA-XS. Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\mathcal{W}_{LoRA-XS}=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
@@
$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies$
$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\subseteq\text{Col}(B).$
@@ B.2 Proof of Lemma 2
The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}.$
@@
$y\frac{\partial X}{\partial R}=yA^{\top}\implies\frac{\partial L}{\partial R}=yA^{\top}=sB^{\top}gA^{\top}\$(17)
$\text{Therefore, \quad}\boxed{g^{R}_{\text{LoRA-XS}}=sB^{\top}gA^{\top}}$ (18)
B.3 Proof of Theorem 3
Theorem.
For full-rank AA and $B$ matrices, the optimal solution for the objective
@@ B.3 Proof of Theorem 3
$\text{Therefore,}\quad\boxed{g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}}$ (24)
B.4 Proof of Theorem 4
@@
Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate. This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$
@@
Next, we note that matrices $B^{\top}B\in\mathbb{R}^{r\times r}$ and $AA^{\top}\in\mathbb{R}^{r\times r}$ are positive definite since they are positive semi-definite and matrices $B$ and $A$ are full-rank (i.e., with rank $r$) matrices, which means that $B^{\top}B$ and $AA^{\top}$ have non-zero eigenvalues. Therefore, $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ are also positive definite, implying that there exist matrices $X$ and $Y$ such that $(B^{\top}B)^{-1}=YY^{\top}$ and $(AA^{\top})^{-1}=XX^{\top}$ (e.g., one can find such matrices using Cholesky decomposition). Then, we have
$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}$ $=\langle g^{R}_{\text{LoRA-XS}},YY^{\top}g^{R}_{\text{LoRA-XS}}XX^{\top}\rangle_{F}$
$=\langle Y^{\top}g^{R}_{\text{LoRA-XS}}X,Y^{\top}g^{R}_{\text{LoRA-XS}}X\rangle_{F}$
$=\|Y^{\top}g^{R}_{\text{LoRA-XS}}X\|_{F}^{2}\geq 0.$
@@ B.5 Proof of Theorem 5
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\quad\text{ but not when }\quad\tilde{g}=sBg^{R}_{LoRA-XS}A.$
@@ B.6 Proof of Theorem 6
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then the update given by LoRA-SB, $\Delta(B_{init}R_{init}A_{init})$ , is the best low-rank approximation of full fine-tuning update, $\Delta W.$
@@
Appendix C Simulating the First Step of Full Fine-Tuning Under AdamW
Our initialization is designed to approximate the first update step that would occur during full fine-tuning using the AdamW optimizer, which is also used in LoRA-SB training. AdamW computes the parameter update using both first and second moment estimates of the gradient. At the first step, these moments are initialized to zero, so the update becomes:
$\theta_{1}=\theta_{0}-\alpha\cdot\frac{g_{1}}{\sqrt{g_{1}^{2}+\epsilon}}\approx-\alpha\cdot\text{sign}(g_{1})$
where $g_{1}$ is the gradient at the first step, $\epsilon$ is a small constant for numerical stability, and $\alpha$ is the learning rate. Due to zero-initialization and bias correction, the direction of the update is approximately the element-wise sign of the gradient.
To simulate this behavior in our low-rank initialization, we use:
$\Delta W_{\text{avg}}=-\eta\cdot\text{sign}\left(\sum_{i=1}^{n}\nabla_{W}\mathcal{L}(W_{0},x_{i})\right)$
This reflects the direction of the first AdamW step averaged over a mini-batch. By using the sign of the gradient sum, we ensure our initialization aligns with the dynamics of AdamW, leading to a consistent and faithful approximation of full fine-tuning updates within the low-rank subspace.
Appendix D Algorithm
We provide a pseudo-code implementation of our method in Algorithm 1.
Algorithm 1 LoRA‑SB, PyTorch‑like
1:def initSB(model, D)
2:   # Estimate gradient with n samples
3:   $\Delta W_{\mathrm{avg}}\leftarrow$ est_grad(model, D, n)
4:   # Initialize B, R, A
5:   $(B,R,A)\leftarrow$ trunc_SVD($\Delta W_{\mathrm{avg}}$)
6:   # Convert to LoRA‑SB model
7:   sb_model $\leftarrow$ lora_SB(model, B, R, A)
8:    return sb_model
9:
10:# Load pre‑trained model
11: model $\leftarrow$ AutoModel(base_model)
12:# Initialize LoRA‑SB with D
13:sb_model $\leftarrow$ initSB(model, D)
14:# Train, only R trainable
15:trainer $\leftarrow$ Trainer(sb_model,…)
16:trainer.train()
Appendix E Optimal Gradient Approximation is Important!
As discussed in Section 4, optimal gradient approximation plays a key role in the effectiveness of LoRA-SB. In Figure 3, we compare the loss curves of models trained with and without this component on Mistral-7B. While both variants begin with similar performance due to effective initialization, LoRA-SB with optimal gradient approximation converges to substantially lower loss values, highlighting its contribution to improved optimization.
Figure 3: Training loss for Mistral-7B, highlighting the impact of optimal gradient approximation.
Appendix F Training Time Overhead vs LoRA-XS
As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset. Table 6 presents the associated training time overhead for these computations, compared to LoRA-XS. The results show that the additional overhead is negligible, adding just 242–4 minutes compared to the total training time of 353–5 hours per epoch ($\approx 1.1\%$ to $1.3\%$). Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training. Notably, the initialization step is highly efficient, as we directly compute the truncated SVD using optimized PyTorch libraries (torch.svd_lowrank). For reference, this computation takes less than one second for each of the entire LLMs used in our experiments.
Model Overhead Training Time/Epoch
Mistral-7B 0:02:01 3:03:57
Gemma-2 9B 0:03:46 4:13:24
Llama-3.2 3B 0:03:54 4:54:31
Appendix G Inference Overhead vs LoRA
LoRA-SB introduces a minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, and the need for higher ranks to achieve comparable performance to LoRA. We benchmark the inference-time FLOPs and MACs across various models and find that the overhead is negligible. This comparison is presented in Table 7, showing that the additional overhead of LoRA-SB is negligible.
@@ Appendix G Inference Overhead vs LoRA
RoBERTa-large LoRA 88 $77.86$ G $155.79$ G
@@
Appendix H Experiment Details
We use PyTorch (33) and the HuggingFace Transformers library (48) for our implementations. We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds. To save memory, we initialize base models in torch.bfloat16 precision. We trained all models using the AdamW optimizer (28). We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples. The samples are randomly selected from the training set in each run.
For arithmetic and commonsense reasoning tasks, we set up Mistral-7B, Gemma-2 9B, and Llama-3.2 3B with hyperparameters and configurations listed in Table 8. We adopted most settings from previous studies (18) but conducted our own learning rate sweep. Following LoRA-XS guidelines, we set $\alpha=r$ for their baseline configuration.
For the GLUE benchmark using RoBERTa-large, you can find the hyperparameter details in Table 9. We mostly adhered to the original configurations from the LoRA paper (17) but adjusted the learning rate through a sweep. In line with LoRA-XS settings, we fixed $\alpha$ at $16$ for their baseline.
For all tasks, we followed the baseline configurations provided in the PiSSA (30), rsLoRA (20), DoRA (26), and LoRA-Pro (46) papers for our comparisons.
@@ Appendix H Experiment Details
Appendix I Dataset Details
The MetaMathQA dataset (50) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information. We assess this dataset using two benchmarks: GSM8K (8), which consists of grade-school math problems requiring multi-step reasoning, and MATH (16), which presents difficult, competition-level math problems. Evaluation focuses solely on the final numeric answer.
CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (18). Each example is framed as a multiple-choice question where the model generates the correct answer without explanations. We use the prompt template from (18). The individual datasets used are described below: